Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
menzione - 16:06 Wednesday 06 July 2022 (56364) Print this report
Operator Report - Morning shift

The planned ISC activity on TCS locking (Pinto, Casanueva, Taranto) started at 07:00 UTC and went on without major problems for the whole shift.
 

Parallel activity:
SQZ - measurements (Ding, De Laurentis, Vahlbruch)
DAQ - 12:05 UTC - All the Fbs servers were restarted.

Automation
At 09:30 UTC DET_MAIN Metatron node got stuck. I put it on Init,  FORCE_CHECK_CLOSE, and then on SHUTTER_CLOSED.

Sub-system reports

DAQ
TCSAgilisRot VPM process crashed, I tried to restart it without success. Experts fixed the problem.
The crash of the TCSAgilisRot process highlighted a pre-existing DAQ problem that was remotely fixed by Rhys Poulton and by Carbognani in TCS Room (replugging the USB port).

DET
At 12:53 UTC the spot of B1p was in the lower part of the screen. I fixed the position via VPM (SDB1_LC):
- Used fixed setpoints (spot again in the center of the screen)
- Open B5 beam drift
- Close B5 beam drift
- (after few minutes) Use floating setpoints
- Open B5 beam drift

TCS
The bug on TCSAgilisRot has been fixed but the optical problem with the rotor that moves a wave plate is still present and must be solved.

Oncall events

DAQ
(06-07-2022 10:00 - 06-07-2022 11:00) Other
Status: Ended
Description: TCSAgilisRot VPM process crashed, I tried to restart it without success. Experts fixed the problem.

Images attached to this report
Search Help
×

Warning

×